OpenStudy (anonymous):

. 1.6x + 2y = 0 4x + y = -1 substitution

OpenStudy (australopithecus):

y = -1-4x 1.6x + 2(-1-4x) = 0 solve for x what ever you get as x sub into one of the original equations to solve for y

OpenStudy (australopithecus):

Problem explained

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow y = -1 - 4x }\) By transposing the 2nd equation, I got this. Now, we'd just replace y with -1 - 4x in the first equation. We get: \(\Large \color{purple}{\rightarrow 1.6x + 2(-1 - 4x) = 0 }\) Can you continue?

OpenStudy (anonymous):

um still consfused

OpenStudy (lgbasallote):

where?

OpenStudy (anonymous):

where do i put in the x?

OpenStudy (lgbasallote):

what do you mean x? do you get the first step? y = -1 - 4x?

OpenStudy (anonymous):

but i mean what is y its all over the place

OpenStudy (anonymous):

and no

Parth (parthkohli):

Can you solve one-variable linear equations?

OpenStudy (lgbasallote):

okay let me explain it...

OpenStudy (lgbasallote):

as you can see you have two variables here right x and y?

OpenStudy (anonymous):

yes and then what

OpenStudy (anonymous):

no i cant solve one linear equations

OpenStudy (lgbasallote):

to make our lives easier, we would like to have only ONE variable. this is where "substitution" comes into play. basically, we "substitute" one of the variables into the other so there would only be one variable...got it?

OpenStudy (anonymous):

uhu